OpenStudy (anonymous):

What is the length of the base of a triangle with height 20 cm and area 150 cm2?

OpenStudy (anonymous):

Area = 1/2 b * h 150 = 1/2 * b * 20 b = 15cm

OpenStudy (anonymous):

For a triangle: Area = .5*base*height For you: Area = 150 and height = 20 Plugging those into the formula, you get 150=.5%2Abase%2A20 Multiply .5 and 20: 150=10%2Abase To get base by itself, divide by 10 on both sides: 15=base The base is 15 cm.